# V2V-GoT: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multimodal Large Language Models and Graph-of-Thoughts

### v3

@@ Abstract
Current state-of-the-art autonomous vehicles could face safety-critical situations when their local sensors are occluded by large nearby objects on the road. Vehicle-to-vehicle (V2V) cooperative autonomous driving has been proposed as a means of addressing this problem, and one recently introduced framework for cooperative autonomous driving has further adopted an approach that incorporates a Multimodal Large Language Model (MLLM) to integrate cooperative perception and planning processes. However, despite the potential benefit of applying graph-of-thoughts reasoning to the MLLM, this idea has not been considered by previous cooperative autonomous driving research. In this paper, we propose a novel graph-of-thoughts framework specifically designed for MLLM-based cooperative autonomous driving. Our graph-of-thoughts includes our proposed novel ideas of occlusion-aware perception and planning-aware prediction. We curate the V2V-GoT-QA dataset and develop the V2V-GoT model for training and testing the cooperative driving graph-of-thoughts. Our experimental results show that our method outperforms other baselines in cooperative perception, prediction, and planning tasks.
@@ I Introduction
In this paper, we propose a new graph-of-thoughts reasoning framework designed for MLLM-based cooperative autonomous driving, as illustrated in Figure 1. We follow the problem setting from V2V-LLM [12], where all CAVs share their perception features with a MLLM. Any CAV can ask the MLLM to answer a driving-related question. The MLLM fuses the perception features from all CAVs and performs inference to provide the answer. Unlike V2V-LLM [12] and DriveLM [9], we propose a novel graph-of-thoughts reasoning framework designed for cooperative autonomous driving scenarios, as shown in Figure 1 and 2.
To verify the effectiveness of our proposed idea, we first curate the training and testing QA data samples of our proposed V2V-GoT-QA dataset, which is built on top of V2V4Real [16] cooperative driving dataset. Our V2V-GoT-QA includes 99 types of QA samples, as shown in Figure 2. Our proposed occlusion-aware perception questions (Q1 - Q4) consider visible, occluding, and invisible objects. Our proposed planning-aware prediction questions (Q5 - Q7) include prediction by perception features and prediction by other CAVs’ current planned future trajectories. Our planning questions (Q8 - Q9) provide the suggested action settings and waypoints of future trajectories to avoid potential collisions. Different types of QA are connected by a directed edge in our graph-of-thoughts, as shown in Figure 1. The answer of the parent QA is used as the input context of the child QA.
In addition to the newly curated dataset, we also develop our baseline model V2V-GoT, as shown in Figure 3. Unlike V2V-LLM [12] and DriveLM [9] that only take the perception features at the current timestep as visual input, our V2V-GoT includes the perception features from both current and previous timesteps. This design can better capture the temporal dynamics of the surrounding driving scenes to provide better prediction and planning performance.
In our experiments, we compare the final planning performance of our proposed V2V-GoT and other baseline methods from the prior work V2V-LLM [12], including no fusion, early fusion, and intermediate fusion. Our experimental results show that our proposed V2V-GoT outperforms all other baselines in the planning tasks, achieving the lowest collision rates and L2 errors between the ground-truth trajectories and model outputs. Furthermore, we also conduct an ablation study to verify that the newly proposed QA types in our graph-of-thoughts, namely occlusion-aware perception QAs (Q1 - Q4) and planning-aware prediction QAs (Q5 - Q7), are helpful in achieving better cooperative perception, prediction, and planning performance for MLLM-based cooperative autonomous driving.
@@ III-A Problem Setting
We build on the MLLM-based cooperative autonomous driving problem setting from V2V-LLM [12]. The MLLM fuses the perception features from all CAVs and performs inference to answer a question. Unlike the prior work V2V-LLM [12], we propose a novel graph-of-thoughts reasoning framework, including the ideas of occlusion-aware perception and planning-aware prediction. We design 99 types of question-answer pairs in our V2V-GoT-QA dataset, including perception, prediction, and planning, as illustrated in Figure 2. Different types of QAs are connected by directed edges in the graph-of-thoughts, as shown in Figure 1 and  4(a). The answer of the parent node QA is used as the input context of the child node QA.
@@ IV-A Architecture
We extend V2V-LLM [12]’s model to build our V2V-GoT model, as shown in Figure 3. LLaVA [25] is used as the base MLLM architecture. Unlike the original LLaVA [25] that uses an image encoder, we apply a LiDAR-based 3D object detector, PointPillars [26], to extract perception features from each individual CAV’s point cloud. Unlike V2V-LLM [12]’s model that only uses the perception features at the current timestep, our model uses the perception features at the current and previous timesteps from all CAVs as the input of the project layers in the MLLM to generate the visual tokens. The MLLM takes the visual tokens and the language tokens from the question and the context as input and generates the final answer in the natural language format.
@@ VIII Conclusion
In this work, we propose a novel graph-of-thoughts reasoning framework for MLLM-based cooperative autonomous driving. Our proposed graph-of-thoughts includes novel reasoning steps, such as occlusion-aware perception and planning-aware prediction, which are designed to take advantage of V2V information sharing in cooperative driving scenarios and the multimodal understanding ability of MLLMs. To verify the effectiveness of our proposed ideas, we curate the V2V-GoT-QA dataset and develop the V2V-GoT model. Our experimental results show that our proposed method outperforms all other baseline methods. Moreover, the ablation study further indicates that our innovative designs in the graph-of-thoughts, such as occlusion-aware perception and planning-ware prediction, improve overall cooperative perception, prediction, and planning performance.

### v4

@@ Abstract
Current state-of-the-art autonomous vehicles could face safety-critical situations when their local sensors are occluded by large nearby objects on the road. Vehicle-to-vehicle (V2V) cooperative autonomous driving has been proposed as a means of addressing this problem, and one recently introduced framework for cooperative autonomous driving has further adopted an approach that incorporates a Multimodal Large Language Model (MLLM) to integrate cooperative perception and planning processes. However, despite the potential benefit of applying graph-of-thoughts reasoning to the MLLM, this idea has not been considered by previous cooperative autonomous driving research. In this paper, we propose a novel graph-of-thoughts framework specifically designed for MLLM-based cooperative autonomous driving. Our graph-of-thoughts includes our proposed novel ideas of occlusion-aware perception and planning-aware prediction. We curate the V2V-GoT-QA dataset and develop the V2V-GoT model for training and testing the cooperative driving graph-of-thoughts. Our experimental results show that our method outperforms other baselines in cooperative perception, prediction, and planning tasks. Our code and dataset are released to facilitate open-source research at https://eddyhkchiu.github.io/v2vgot.github.io/.
@@ I Introduction
In this paper, we propose a new graph-of-thoughts reasoning framework designed for MLLM-based cooperative autonomous driving, as illustrated in Figure 1. We follow the problem setting from V2V-LLM [2], where all CAVs share their perception features with a MLLM. Any CAV can ask the MLLM to answer a driving question. The MLLM fuses the perception features from all CAVs and performs inference to provide the answer. Unlike previous work [2, 16], we propose a novel graph-of-thoughts reasoning framework specifically designed for cooperative autonomous driving scenarios, as illustrated in Figure 1 and 2.
To verify the effectiveness of our proposed idea, we first curate the V2V-GoT-QA dataset, built upon the V2V4Real [24] cooperative driving dataset. Our V2V-GoT-QA includes $9$ different types of QA samples, as illustrated in Figure 2. Our proposed occlusion-aware perception questions (Q1 - Q4) consider visible, occluding, and invisible objects. Our proposed planning-aware prediction questions (Q5 - Q7) include prediction by perception features and prediction by other CAVs’ current planned future trajectories. Our planning questions (Q8 - Q9) provide the suggested action settings and suggested waypoints of future trajectories to avoid potential collisions. Different types of QA are connected by a directed edge in our graph-of-thoughts, as shown in Figure 1. The answer of the parent node QA is used as the input context of the child node QA.
In addition to the newly curated dataset, we also develop our baseline model V2V-GoT, as shown in Figure 3. Unlike the prior work V2V-LLM [2] and DriveLM [16] that only take the perception features at the current timestep as visual input, our V2V-GoT includes the perception features from both current and previous timesteps to better capture the temporal dynamics of the surrounding objects.
In our experiments, we compare the final planning performance of our proposed V2V-GoT and other baseline methods from the prior work V2V-LLM [2], including no fusion, early fusion, and intermediate fusion. Experimental results show that our model outperforms all other baselines in the planning tasks, achieving the lowest collision rates and L2 errors. Furthermore, our ablation study also verifies that the newly proposed QA types in our graph-of-thoughts, namely occlusion-aware perception QAs (Q1 - Q4) and planning-aware prediction QAs (Q5 - Q7), are helpful in achieving better cooperative perception, prediction, and planning performance for MLLM-based cooperative autonomous driving.
@@ III-A Problem Setting
We build on the MLLM-based cooperative autonomous driving problem setting from V2V-LLM [2]. The MLLM fuses the perception features from all CAVs and performs inference to answer a question. Unlike the prior work V2V-LLM [2], we propose a novel graph-of-thoughts reasoning framework, including the ideas of occlusion-aware perception and planning-aware prediction. We design $9$ types of question-answer pairs in our V2V-GoT-QA dataset, including perception, prediction, and planning, as illustrated in Figure 2. Different types of QAs are connected by directed edges in the graph-of-thoughts, as shown in Figure 1 and  4(a). The answer of the parent node QA is used as the input context of the child node QA.
@@ IV-A Architecture
We build our V2V-GoT model with LLaVA-v1.5-7b [11] as the base MLLM architecture, as shown in Figure 3. Unlike the original LLaVA [11] that uses an image encoder, we apply a LiDAR-based 3D object detector, PointPillars [9], to extract perception features from each individual CAV’s point cloud. Unlike V2V-LLM [2]’s model that only uses the perception features at the current timestep, our model uses the perception features at the current and previous timesteps from all CAVs as the input of the project layers in the MLLM to generate the visual tokens. The MLLM takes the visual tokens and the language tokens from the question and the context as input and generates the final answer in the natural language format.
@@ VIII Conclusion
In this work, we propose a novel graph-of-thoughts reasoning framework for MLLM-based cooperative autonomous driving. Our proposed graph-of-thoughts includes occlusion-aware perception and planning-aware prediction, which are designed to take advantage of V2V information sharing in cooperative driving scenarios and the multimodal understanding ability of MLLMs. To verify the effectiveness of our proposed ideas, we curate the V2V-GoT-QA dataset and develop the V2V-GoT model. Our experimental results show that our proposed method outperforms all other baseline methods. Moreover, the ablation study further indicates that our innovative designs in the graph-of-thoughts improve overall cooperative perception, prediction, and planning performance. To facilitate open-source research, we have publicly released our V2V-GoT-QA dataset and V2V-GoT code.